OpenStudy (anonymous):

writing piecewise functions? f(x)=abs(2x)-abs(x-2)+3

OpenStudy (anonymous):

\[f(x)=\left| 2x \right|-\left| x-2 \right|+3\]

OpenStudy (paxpolaris):

this function can be broken into 3 pieces

OpenStudy (anonymous):

i know that much...and dont you write one as it is and the other with the negative sign?

OpenStudy (paxpolaris):

- when \(\large x<0\) ... both (2x) and(x-2) are negative - when \(\large 0\le x < 2\) ... (x-2) is negative but (2x) is not - and when \(\large 2\le x\) ... both (2x) and(x-2) are not negative

OpenStudy (anonymous):

oh okay and how did you find which one is negative and which one isnt?

OpenStudy (paxpolaris):

if 2x is negative \[\large \implies 2x < 0 \]\[\large \implies x < 0 \] if (x-2) is negative \[\large \implies x-2 < 0 \]\[\large \implies x-2+2 < 0+2 \]\[\large \implies x < 2 \]

OpenStudy (paxpolaris):

whenever something inside the absolute value is negative you flip its sign.

OpenStudy (anonymous):

oh i understand now thank you:)

OpenStudy (paxpolaris):

ok.. thanks for the medal